This blog follows the progress of the LRO mission through Integration and Testing at NASA's Goddard Space Flight Center and launch site processing at KSC\Astrotech. Its purpose is to enable communication to the entire LRO Team about the status of ongoing activities.

LRO was launched June 18th, 2009 at 5:32pm EDT from KSC. This BLOG will follow the progress of the mission as LRO travels to the Moon and establishes orbit around it.

This BLOG will be periodically updated during LRO's early mission but as the nominal mission unfolds the official NASA LRO website and the LRO Science Instrument's own websites will evolve into a more interesting sources of LRO science results as postings appear there first and LRO engineering and operations (source of this site) will become increasingly routine.




Saturday, November 8, 2008

November 8, 2008 @3:30 PM
We have completed the following: Installation of the LVA cold ring,
installation of the Star Tracker Stimulus with heaters and thermocouples, installation LOLA shower head with heaters and thermocouples, but without the fiber optic cables; they will be installed on Monday, additional thermocouples were installed on the propulsion system cat bed heaters.
At this point we are still working on modifications to the Solar Array Gimbal thermal hardware; we are removing some blankets, installing Kapton tape on the harness and adding some heaters to the gimbal assembly. This activity will consume the remainder of the day. Other blanket work is going on in parallel to this task.
There is no work scheduled for tomorrow, Sunday 11/9/2008; we will resume chamber break activities on Monday morning, 11/9/2008.

Status 11/8 @9:30am

Completed the installation of the LVA cold ring. Currently we are installing the blanket over the test connector panels, TCs to the catbed heaters and the Star Tracker OGSE.

The SAS power harness has been secured to the enclosure by Brian.

Other things on the docket today include installing the LOLA GSE and opening up the blanket on the SAS gimbal and installing TCs.

Friday, November 7, 2008

Status 11/7 @10:30pm

We have completed the installation of the test harnesses, including the RF cables and the battery GSE rack.

Start of the LVA cold ring installation is 7am Saturday morning.

Status 11/7 @ 8:30pm

So far we have managed to open the chamber door, perform the radiation test, remove the radiation source, remove the enclosure panels, install the catwalks, certify the battery GSE rack, open blankets around the LVA and test connector panel; and install some of the thruster covers.

Still need to remove the test connector panel cover, test the pyro valve NSIs, safe to mate the battery GSE interface and install the test connectors.

Plan for Saturday.

Here is the plan for Saturday.

- NO POWERED WORK.

We can’t work on the SAS gimbal and do testing at the same time. We will try to resume powered testing on Monday
- Install lower thruster covers. 7a
- Install LVA cold ring. 7a
- Install RF cable. 9a
- Install Star Tracker GSE. 8:30a
- Install LOLA GSE. 9a
- Install heaters & TCs on GSE. (after GSE installed)
- Work on SAS gimbal harness. 7a (all day)
- Install TCs near catbed heaters. (after cold ring installed)

Current Plans for tonight.

We started to backfill the chamber at 1:00 today. We should be ready to open the door around 4pm.

- 4pm -5p: Radiation Safety enters chamber to remove source/take sample & let us know we are clear to enter the chamber.

- 5p: Chamber & Blanket folks go in to remove enclosure panels.

- Contamination removes witness plates

- Blanket folks make access for installation of the LVA cold ring and access to the Connector Test Panel.

- Mechanical and /or Chamber folks install +X and –X cat walks. (around 7p???)

- Mechanical to remove Test connector panel cover.

- Electrical folks to help prop test NSIs

- Electrical folks to configure test connector panel.

- Battery GSE STM.

- Chamber folks to remove TCs at –X end.

- SAS Gimbal blanket mods start

Door opening now estimated to be around 3pm.

Waiting for the payload table (the platform the spacecraft is sitting on) to get warmer. Then about 2 hrs to backfill chamber.

Waiting to open the door.

Looks like we won't be able to open the chamber door until around lunch time. Could be as late as 1p-2p. :-(